# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ 5.1 Intuition behind the algorithm
12:         if hedgeQuery then
13:              Query mentor and observe $\pi^{m}(x_{t})$
14:              $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ for all $\pi\in\tilde{\Pi}$
15:              $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
16:              $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
17:              $\pi_{t}\leftarrow\operatorname*{arg\,min}_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
18:         else
19:              $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
20:              Sample $\pi_{t}\sim P$
21:         if $\min_{(x,y)\in X:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then $\triangleright$ Ask for help if out-of-distribution
22:              Query mentor and observe $\pi^{m}(x_{t})$
23:              $X\leftarrow X\cup\{(x_{t},\pi^{m}(x_{t}))\}$
24:         $\triangleright$ Otherwise, follow Hedge’s chosen policy
25:              Take action $\pi_{t}(x_{t})$
@@
We prove the following theorem parametrized by $\varepsilon$:
@@
Let $\mathcal{Y}=\{0,1\}$. Assume $\pi^{m}\in\Pi$ where either (1) $\Pi$ has finite VC dimension $d$, $\boldsymbol{x}$ is $\sigma$-smooth, and $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ or (2) $\Pi$ has finite Littlestone dimension $d$. Then for any $T\in\mathbb{N}$ and $\varepsilon>0$, Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(1/\varepsilon)\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
In Case 1, the expectation is over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm. Also, $R_{T}$ and $Q_{T}$ clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate bounds.
To obtain subconstant regret and sublinear queries, we can choose $\varepsilon=T^{\frac{-2n}{2n+1}}$. This also satisfies the requirement of $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$ for large enough $T$.
@@
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
Although our focus is the product of payoffs, Algorithm 1 also guarantees subconstant additive regret:
Theorem 5.3.
$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
@@ 5.2 Proof sketch
The formal proof of Theorem 5.1 can be found in Appendix B, but we outline the key elements here. The regret analysis consists of two ingredients: analyzing the Hedge component, and analyzing the “ask for help when out-of-distrubtion” component. The former will bound the number of mistakes made by the algorithm (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s), and the latter will bound the cost of any single mistake. We must also carefully show that the latter does not result in excessively many queries, which we do via a novel packing argument.
We begin by formalizing two notion of approximating a policy class:
@@
Let $U$ be the uniform distribution over $\mathcal{X}$. For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon$.
@@
A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\boldsymbol{x}\in\mathcal{X}^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(x_{t})=\tilde{\pi}(x_{t})$ for all $t\in[T]$.
@@
Lemma 5.1 (Lemma 7.3.2 in Haghtalab (2018) [^13] See also Haussler & Long (1995) or Lemma 13.6 in Boucheron et al. (2013) for variants which are less convenient for our purposes.).
@@
An adversarial cover is a perfect cover by definition. The following lemma establishes that a smooth $\varepsilon$-cover is a good approximation for any sequence of $\sigma$-smooth distributions.
Lemma 5.3 (Equation 2 and Lemma 3.3 in Haghtalab et al. (2024)).
Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions. If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then $\operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(x_{t})\neq\tilde{\pi}(x_{t}))\right]\in O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$.
@@
We cannot afford to query on every time step, however. Recently, Russo et al. (2024) analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”. They proved a regret bound when a query is issued on each time step with fixed probability $p$. Lemma 5.4 restates their result in a form that is more convenient for us (see Appendix B for details). Although their result is stated for non-adaptive adversaries, we explain in Section B.3 how their argument easily generalizes to adaptive adversaries. Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
Lemma 5.4 (Lemma 3.5 in Russo et al. (2024)).
Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p$, HedgeWithQueries enjoys the regret bound
$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\ell(t,\pi)\leq\frac{2\log|\tilde{\Pi}|}{p^{2}}$
@@
We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case). This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
The other key ingredient of the proof is analyzing the “ask for help when out-of-distribution” component. Combined with the local generalization assumption, this allows us to fairly easily bound the cost of a single mistake (Lemma B.2). The trickier part is bounding the number of resulting queries. It is tempting to claim that the inputs queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true. Instead, we provide a novel method for bounding the number of data points (i.e., queries) needed to cover a set with respect to the realized actions of the algorithm (Lemma B.7). This is in contrast to vanilla packing arguments which consider all data points in aggregate. Our method may be useful in other contexts where a more refined packing argument is needed.
@@ 6 Conclusion and future work
There remain some technical questions within this paper’s model. One question is whether the time complexity of Algorithm 1 be improved, which currently stands at $\Omega(|\tilde{\Pi}|\cdot T)$ plus the time to compute the $\varepsilon$-cover. Also, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension).
We are also interested in alternatives to the local generalization assumption. We should expect some assumption to be necessary: if not, the payoff function $\mu(x,y)=\mathbf{1}(\pi^{m}(x)=y)$ means that the agent must make zero mistakes (with high probability), which turns out to be impossible even for $\sigma$-smooth $\boldsymbol{x}$ and finite VC dimension (Theorem E.3). One possible alternative is Bayesian inference. We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
Finally, we are excited to apply the ideas in this paper to Markov Decision Processes (MDPs): specifically, MDPs where some actions are irreversible (“non-communicating”) and the agent only gets one attempt (“single-episode”). In such MDPs, the agent must not only avoid catastrophe but also obtain high reward. As discussed in Section 2, very little theory exists for RL in non-communicating single-episode MDPs. Can an agent learn near-optimal behavior in high-stakes environments while becoming self-sufficient over time? Formally, we pose the following open problem:

### v5

@@ 5.2 Local generalization
[…]
     if hedgeQuery [^9] The reader may notice that we do not update $S$ in this case. This is simply because those updates are not necessary for the desired bounds and omitting these updates simplifies the analysis. then
        Query mentor and observe $\pi^{m}(x_{t})$
        $\ell(t,\pi)\leftarrow\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ for all $\pi\in\tilde{\Pi}$
        $\ell^{*}\leftarrow\min_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
        $w(\pi)\leftarrow w(\pi)\cdot\exp(-\eta(\ell(t,\pi)-\ell^{*}))$ for all $\pi\in\tilde{\Pi}$
        $\pi_{t}\leftarrow\operatorname*{arg\,min}_{\pi\in\tilde{\Pi}}\ell(t,\pi)$
     else
        $P(\pi)\leftarrow w(\pi)/\sum_{\pi^{\prime}\in\tilde{\Pi}}w(\pi^{\prime})$ for all $\pi\in\tilde{\Pi}$
        Sample $\pi_{t}\sim P$
     if $S=\emptyset$ or $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||>\varepsilon^{1/n}$ then
        Ask for help if out-of-distribution
        Query mentor (if not already queried this round) and observe $\pi^{m}(x_{t})$
        $S\leftarrow S\cup\{(x_{t},\pi^{m}(x_{t}))\}$
     else
        Otherwise, follow Hedge’s chosen policy
        Take action $\pi_{t}(x_{t})$
@@
Conceptually, the algorithm only needs to be able to detect when an input is unfamiliar. While this task remains far from trivial, we argue that it is more tractable than fully constructing a suitable encoding. See Section 6 for a discussion of potential future work on this topic.
We note that these encoding-related questions apply similarly to the more standard assumption of Lipschitz continuity. In fact, Lipschitz continuity implies local generalization when the mentor is optimal (Proposition E.1). We also mention that without local generalization, avoiding catastrophe is impossible even when the mentor policy class has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth (Theorem E.2).
5.3 Main result
For simplicity, here we only state our results for $\mathcal{Y}=\{0,1\}$; Appendix C extends our result to many actions using the standard “one versus rest” reduction. We first prove regret and query bounds parametrized by $\varepsilon$:
@@ 5.3 Main result
[…]
$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{dL}{\sigma}T\varepsilon^{1+1/n}\log(T+1/\varepsilon)\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\!\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T\!+\!1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
In Case 1, the expectation is over the randomness of both $\boldsymbol{x}$ and the algorithm, while in Case 2, the expectation is over only the randomness of the algorithm. The bounds clearly have no dependence on $\sigma$ in Case 2, but we include $\sigma$ anyway to avoid writing two separate sets of bounds.
To obtain subconstant regret and sublinear queries, we can choose $\varepsilon=T^{\frac{-2n}{2n+1}}$. This also satisfies $2L\varepsilon^{1/n}\leq\mu_{0}^{m}$ for large enough $T$.
@@
$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)\right)$
Before proceeding to the proof sketch, we highlight some advantages of our algorithm.
Limited knowledge required. Our algorithm needs to know $\Pi$, as is standard. However, the algorithm does not need to know $\sigma$ (in the smooth case) or $L$. Although Algorithm 1 as written does require $T$ as an input, it can be converted into an infinite horizon/anytime algorithm via the standard “doubling trick” (see, e.g., Slivkins et al., 2019).
Unbounded environment. Our algorithm can handle an unbounded input space: the number of queries simply scales with the maximum distance between observed inputs in the form of $\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]$.
Simultaneous bounds for all $\boldsymbol{\mu}$. Recall that the agent never observes payoffs and only learns from mentor queries. This means that the agent’s behavior does not depend on $\boldsymbol{\mu}$ at all. In other words, the distribution of $(\boldsymbol{x},\boldsymbol{y})$ depends on $\pi^{m}$ but not $\boldsymbol{\mu}$. Consequently, a given $\pi^{m}$ induces a single distribution $(\boldsymbol{x},\boldsymbol{y})$ which satisfies the bounds in Theorem 5.2 simultaneously for all $\boldsymbol{\mu}$ satisfying local generalization.
@@ 5.4 Proof sketch
The formal proof of Theorem 5.1 can be found in Appendix B, but we outline the key elements here. The regret analysis consists of two ingredients: analyzing the Hedge component and analyzing the “ask for help when out-of-distribution” component. The former will bound the number of mistakes made by the algorithm (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s), and the latter will bound the cost of any single mistake. We must also show that the latter does not result in excessively many queries, which we do via a novel packing argument.
We begin by formalizing two notions of approximating a policy class:
@@
Let $U$ be the uniform distribution over $\mathcal{X}$. For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ if for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon$.
@@
A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ if for every $\boldsymbol{x}\in\mathcal{X}^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(x_{t})=\tilde{\pi}(x_{t})$ for all $t\in[T]$.
An adversarial cover is a perfect cover by definition. The idea of a smooth $\varepsilon$-cover is that if the probability of disagreement over the uniform distribution is small, then the probability of disagreement over a $\sigma$-smooth distribution cannot be too much larger.
Lemma 5.1.
Let $\tilde{\Pi}$ be a smooth $\varepsilon$-cover of $\Pi$ and let $\mathcal{D}$ be a $\sigma$-smooth distribution. Then for any $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim\mathcal{D}}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon/\sigma$.
Proof.
Define $S(\tilde{\pi})=\{x\in\mathcal{X}:\pi(x)\neq\tilde{\pi}(x)\}$. By the definition of a smooth $\varepsilon$-cover, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[x\in S(\tilde{\pi})]\leq\varepsilon$. Since $\mathcal{D}$ is $\sigma$-smooth, $\Pr_{x\sim\mathcal{D}}[\pi(x)\neq\tilde{\pi}(x)]=\Pr_{x\sim\mathcal{D}}[x\in S(\tilde{\pi})]\leq\\ \Pr_{x\sim U}[x\in S(\tilde{\pi})]/\sigma\leq\varepsilon/\sigma$, as claimed. ∎
@@
Lemma 5.2 (Lemma 7.3.2 in Haghtalab (2018) [^10] See also Haussler & Long (1995) or Lemma 13.6 in Boucheron et al. (2013) for variants of this lemma.).
@@
We cannot afford to query on every time step, however. Recently, Russo et al. (2024) analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”. They proved a regret bound when a query is issued on each time step with fixed probability $p$. Lemma 5.4 restates their result in a form that is more convenient for us. See Appendices B.1 and B.3 for details on our usage of results from Russo et al. (2024). Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
Lemma 5.4 (Lemma 3.5 in Russo et al., 2024).
Assume $\tilde{\Pi}$ is finite. Then for any loss function $\ell:[T]\times\tilde{\Pi}\to[0,1]$ and query probability $p>0$, HedgeWithQueries enjoys the regret bound
$\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi_{t})]-\min_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\operatorname*{\mathbb{E}}[\ell(t,\pi)]\leq\frac{\log|\tilde{\Pi}|}{p^{2}}$
@@
We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ and combine this with Lemmas 5.2 and 5.1 (in the $\sigma$-smooth case) and with Lemma 5.3 (in the adversarial case). This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
The other key ingredient of the proof is analyzing the “ask for help when out-of-distribution” component. Combined with the local generalization assumption, this allows us to fairly easily bound the cost of a single mistake (Lemma B.2). The trickier part is bounding the number of resulting queries. It is tempting to claim that the inputs queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually false.
Instead, we bound the number of data points (i.e., queries) needed to cover a set with respect to the realized actions of the algorithm (Lemma B.6). This contrasts with vanilla packing arguments which consider all data points in aggregate. The key to our analysis is that the number of mistakes made by the algorithm – which we already bounded in Lemma B.1 – gives us crucial information about how data points are distributed with respect to the actions of the algorithm. Our technique might be useful in other contexts where a more refined packing argument is needed and a bound on the number of mistakes already exists.
@@ 6 Conclusion and future work
Remaining technical questions. First, we have not resolved whether our problem is tractable for finite VC dimension and fully adversarial inputs (although Appendix D shows that the problem is tractable for at least some classes with finite VC but infinite Littlestone dimension). Second, the time complexity of Algorithm 1 currently stands at a hefty $\Omega(|\tilde{\Pi}|)$ per time step plus the time to compute $\tilde{\Pi}$. In the standard online learning setting, Block et al. (2022) and Haghtalab et al. (2022) show how to replace discretization approaches like ours with oracle-efficient approaches, where a small number of calls to an optimization oracle are made per round. We are optimistic about leveraging such techniques to obtain efficient algorithms in our setting.
Local generalization. Our algorithm crucially relies on the ability to detect when an input is unfamiliar, i.e., differs significantly from prior observations in a metric space which satisfies local generalization. Without this ability, the practicality of our algorithm would be fundamentally limited. One option is to use out-of-distribution (OOD) detection, which is conceptually similar and well-studied (see Yang et al., 2024 for a survey). However, it is an open question whether standard OOD detection methods are measuring distance in a metric space which satisfies local generalization.
We are also interested in alternatives to local generalization. Theorem E.2 shows that our positive result breaks down if local generalization is removed, so some sort of assumption is necessary. One possible alternative is Bayesian inference. We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
MDPs. Finally, we are excited to apply the ideas in this paper to Markov Decision Processes (MDPs): specifically, MDPs where some actions are irreversible (“non-communicating”) and the agent only gets one attempt (“single-episode”). In such MDPs, the agent must not only avoid catastrophe but also obtain high reward. As discussed in Section 2, very little theory exists for RL in non-communicating single-episode MDPs. Can an agent learn near-optimal behavior in high-stakes environments while becoming self-sufficient over time? Formally, we pose the following open problem:
@@
Impact statement
As AI systems become increasingly powerful, we believe that the safety guarantees of such systems should become commensurately robust. Irreversible costs are especially worrisome, and we hope that our work plays a small part in mitigating such risks. We do not believe that our work has any concrete potential risks that should be highlighted here.
Author contributions
B. Plaut conceived the project. B. Plaut designed the mathematical model with feedback from H. Zhu and S. Russell. B. Plaut proved all of the results. H. Zhu participated in proof brainstorming and designed counterexamples for several early conjectures. B. Plaut wrote the paper, with feedback from H. Zhu and S. Russell. S. Russell supervised the project and secured funding.
Acknowledgements
This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley. This paper also benefited from discussions with many other researchers. We would like to especially thank (in alphabetical order) Aly Lidayan, Bhaskar Mishra, Cameron Allen, Juan Liévano-Karim, Matteo Russo, Michael Cohen, Nika Haghtalab, and Scott Emmons. We would also like to thank our anonymous reviewers for helpful feedback.
@@ A.1 Proof notation
  2. 2.
    Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried. Since each query appears in exactly one section (because each input appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|$.
  3. 3.
    For each $j\in J_{\neg Q}$, let $y^{j}$ be the most frequent action among time steps in $M_{j}$: $y^{j}=\operatorname*{arg\,max}_{y\in\{0,1\}}|\{t\in M_{j}:y=y_{t}\}|$.
  4. 4.
    Let $J_{\neg Q}^{\prime}=\{j\in J_{\neg Q}:a_{j}\neq y^{j}\}$ be the set of sections where the more frequent action is wrong according to $\mu_{f,\boldsymbol{a}}$.
  5. 5.
    Let $M_{j}^{\prime}=\{t\in M_{j}:y_{t}\neq a_{j}\}$ be the set of time steps where the agent chooses the wrong action according to $\mu_{f,\boldsymbol{a}}$, and $x_{t}$ is close to the midpoint of section $j$.
A.2 Proof roadmap